OpenStudy (trisarahtops):

Answer Check!! Find the velocity, v(t), for an object moving along the x-axis if the acceleration, a(t), is a(t) = 2t + sin(t) and v(0) = 4. v(t) = t2 + cos(t) + 3 v(t) = 2 + cos(t) + 1 v(t) = t2 − cos(t) + 5 <--- v(t) = t2 + sin(t) + 4

OpenStudy (sweetburger):

I think to get the velocity equation from the acceleration you integrate. v(t) =t^2 -cos(t) + C 4= 0^2 -cos(0) + C C = 5 looks like you picked the correct answer.

OpenStudy (trisarahtops):

alright thanks for the help :)

OpenStudy (sweetburger):

All you! Gj.

OpenStudy (trisarahtops):

Thanks